A wave is modeled with the function y=1/2 sin3x. Describe the graph of this function, including its period, amplitude, and points of intersection with the x axis.
I have the amp. is 1/2
the period is (2 pi)/3
but don't know how to get the points of intersection. I know one is (0,0)

Answer by lwsshak3(11628)

A wave is modeled with the function y=1/2 sin3x. Describe the graph of this function, including its period, amplitude, and points of intersection with the x axis.
I have the amp. is 1/2
the period is (2 pi)/3
but don't know how to get the points of intersection. I know one is (0,0)
***
Basic equation for sin function: y=ASin(Bx-C), A=amplitude, Period=2π/B, phase shift=C/B.
For given equation: y=1/2 sin3x
Amplitude=1/2
B=3
Period=2π/B=2π/3
C=0
phase shift: none
Because there is no phase shift, the graph is like the basic sin curve with amplitude=1/2 and period=2π/3
As with the basic sin curve, it intersects the x-axis at zero, half the period, and at the end of the period. So, for given function, points of intersection for one period are: (0,0), (π/3,0) and (2π/3,0)